# Do global forecasting models require frequent retraining?

### v1

@@ 4 Results and discussion
In general, the models tested performed better, in absolute terms, on the M5 dataset compared to the VN1 (see the RMSSE and MQL tables in the supplementary materials of the article). This may be due to different factors like the dataset size, the frequency of the time series, the availability of external regressors to account for promotions, events, etc, and the absence of reference hyperparameter values to use during the training process.
Figure 1 shows the forecast accuracy of each model along the different retraining scenarios for the M5 and the VN1 datasets. To simplify comparisons, for each dataset we report the results in relative terms with respect to the corresponding baseline scenario, that is $r=7$ for M5 and $r=1$ for VN1. With the only exception of the CatBoost model, the RMSSE profiles are very stable over the retraining frequencies. Indeed, the accuracy remains practically the same in the M5 dataset and even improves in the VN1 dataset, regardless of the retraining scenario considered. Especially for low periodic retraining scenarios, the performances of most global models are indistinguishable from the baseline, and even if some deterioration is present for higher retraining frequencies, this is less than 5% also for the no retraining setup. These results imply that less frequent retraining does not harm the point forecast accuracy of global models. This can be explained by the fact that, if the data remains stable without significant trends or concept drifts, as is the case for both the M5 and VN1 datasets, the forecasts will accurately track demand over extended periods.
Figure 2 statistically confirms the above results for the M5 dataset 777 The Friedman-Nemenyi test is usually adopted to compare and rank the performance of different models. Here we use it to compare the accuracy produced by the same model over different retraining scenarios. . It is clear that the retraining scenarios are not statistically different in terms of point forecast accuracy at the 5% level. Indeed, even though the mean rank of the continuous retraining scenario is lower, the intervals intersect with the other scenarios, hence we can not differentiate between them. For the VN1 dataset (see supplementary materials), less frequent retrainings are even statistically significant, meaning that periodic retraining does improve accuracy over continuous retraining.
@@
In a similar fashion to Figure 1, Figure 3 summarizes the relative accuracy in a probabilistic forecasting setting, as defined in 3.5. In this context, we observe that for the M5 dataset the accuracy (as measured by Multi Quantile Loss) is clearly an increasing function of the retrainig scenario, meaning that less frequent updates harm the probabilistic forecasting performance of the models. This is true regardless of the method used. Nonetheless, the differences in accuracy are irrelevant for low retraining scenarios and slightly more pronounced for higher retraining levels, but in any case less than 5-6 percentage points. We also note a small difference in the performances of machine learning and deep learning models, where the formers perform consistently better as the retraining scenario increases. For the VN1 dataset, instead, we observe an almost convex relationship between the accuracy and the retraining period. On average, models’ performance improves for low retraining scenarios and then starts deteriorating around $r=4$. The only exception are NBEATSx and NHITS models, designed also for long-term forecasting, which are consistently better compared to the others for longer retraining scenarios, both for point and probabilistic forecasting.
@@
The statistical test to compare probabilistic forecasting performances among the different retraining scenarios on the M5 dataset as depicted in Figure 4 confirms that there are significant differences in accuracy. In particular, lower restraining scenarios tend to produce more accurate probabilistic forecasts, that is, less frequent retraining effectively harms the accuracy. However, this is not true for low retraining scenarios where the differences are not statistically significant, implying that even in the context of probabilistic forecasting the frequency of retaining may be reduced to at least once a month (or every two months). For the VN1 (results in supplementary materials), instead, some level of periodic retaining is better than continuous retraining most of the time.
@@
Figure 3: MQL results for each method and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
Figure 4: M5 Friedman-Nemenyi test results based on MQL.
@@
Figure 5 shows the relative computational time of each retraining scenario for the two datasets. We observe that CT decreases exponentially as the retraining scenario increases. On average, reductions are similar for both, the M5 and the VN1: going from the baseline to the lowest periodic retraining scenario ($r=14$ or $r=2$) almost halves CT, retraining the models every month often reduces computing time by 75%, and this reduction reaches 90% in the no retraining scenario, where the models are trained just once. However, we observe that there is a strong difference between machine learning and deep learning models in the M5 data. While the former continuously undergoes CT reductions, the latter seems to plateau at 50%, experiencing much lower gains over increasing retraining scenarios. As expected, this difference is less evident in the VN1 dataset, since it is smaller and of lower frequency compared to the M5. These results will have direct implications in terms of costs of forecasting, which in turn may guide in the choice of the model to adopt.
@@
Overall, the results of Figures 1, 3, and 5, combined with the Friedman-Nemenyi tests, suggest that retraining the global models less frequently does not harm (or even improves) forecast accuracy, while at the same time significantly reduces the computing time of producing the forecasts. Thus, extending the retraining time from the standard practice of continuous retaining to some level of periodic retraining allows to effectively manage the computing time, and in turn the costs of forecasting. Indeed, as already observed, the computational time can be directly translated into actual costs for the company. Following Nikolopoulos & Petropoulos (2018) and Fotios Petropoulos & Spiliotis (2024), we assumed some standard costs for computing services to estimate the costs of forecasting associated with each retrain scenario. Note that costs and savings are normalized by the number of SKUs in each datasets so that they can be directly compared and conclusions can be drawn in terms of the frequency of the time series. Figures 6, and 7 show the costs and associated savings for a large retailer, given a fixed computing service cost of \$3.5/hour, 200,000 unique SKUs and 5,000 stores, which is approximately the size of the forecasting problem of Walmart (Spiliotis & Petropoulos, 2024). As expected, the costs decrease exponentially with less frequent retraining. For daily data, forecasting with machine learning models usually costs less than with deep learning models. On average, the continuous retraining scenario costs approximately \$750,000 and this cost drops down to almost \$250,000 in the no retraining scenario, implying direct savings of more than 60%. Moreover, machine learning models permit to reach higher savings. Indeed, even if the models have to be updated, going from retraining every week to retraining every month allows to get direct savings of almost 75%, while it can be less than 30% for deep learning models. This implies that, machine learning methods are a marginally better choice to reduce the costs of forecasting when coupled with less frequent model retraining as the frequency of the data increases. For weekly data, these differences are less evident, meaning that machine learning and deep learning models have usually very similar costs and savings profiles (without considering the Random Forest model, which is usually more than 10 times slower than other methods). This is to be expected since, lower frequencies imply also smaller datasets and lower computing time. In this case, the average cost of forecasting under a continuous retraining scenario is \$250,000 (1/3 of that of daily data) and it drops to \$15,000 in the least frequent retraining scenario, producing direct savings of 90%. Nevertheless, for weekly data too, moving from retraining a model every week ($r=1$) to just once a month ($r=4$) allows to reduce the costs of almost 75%.
@@
It may be argued that for a large retailer (like Walmart) these costs (and savings) may be negligible. However, it has to be noted that the cost reduction obtained from less frequent retraining comes with no shortcomings in terms of forecasting accuracy (especially for point forecasting). Indeed, as already mentioned, models under a periodic retraining scenario result in at least the same (if not even better) forecasting performance compared to the usual practice of continuous retraining. Moreover, periodic retraining may be a good practice even when probabilistic forecasts are needed, balancing costs and accuracy effectively.
@@ 5 Conclusions
In this study, we went beyond the traditional evaluation of forecasting models by also exploring the computational cost associated with generating forecasts. We examined the effects of retraining frequency on the accuracy and computational efficiency of global time series forecasting models. By systematically evaluating various retraining scenarios, ranging from continuous retraining to no retraining at all, across multiple machine learning and deep learning models, we aimed to address a key question in forecasting: is frequent retraining necessary for maintaining high predictive performance? Our findings, derived from an extensive evaluation of ten different global models across two large-scale, real-world retail datasets, challenge the conventional wisdom of continuous model retraining. Indeed, our results indicate that the conventional practice of continuous retraining may not always be justified, and that periodic retraining strategies can offer substantial benefits in terms of computational efficiency without a significant loss in accuracy.
@@
Our empirical evaluation provides evidence that the forecasting accuracy of global models remains stable even when retraining frequency is significantly reduced. For point forecasting, the root mean squared scaled error (RMSSE) results indicate that models trained less frequently do not exhibit substantial performance degradation. In fact, in some cases, particularly for periodic retraining frequencies, we observe marginal improvements in accuracy. For probabilistic forecasting, the results are slightly more nuanced. While less frequent retraining does lead to minor reductions in accuracy as measured by the Multi-Quantile Loss (MQL), the degradation is relatively small (typically within 5-6%). This implies that for most practical applications, especially those where forecasting costs are key considerations, periodic retraining provides a favorable balance between accuracy and efficiency. This suggests that global models, that learn shared dynamics across multiple time series, are robust to the absence of frequent updates, especially in contexts with relatively stable demand patterns.
@@
Considering periodic retraining scenarios allows to effectively manage the computing time, and in turn the costs of forecasting. Indeed, we have shown how the computational time can be directly translated into actual costs for the company. One of the most significant findings of this study is the exponential reduction in computational costs as retraining frequency decreases. The computational time (CT) analysis demonstrates that moving from continuous retraining to a monthly retraining schedule can reduce computational costs by approximately 75%. In the extreme case of no retraining, the cost reductions approach 90%, representing a resource-saving opportunity with no-shortcomings. Furthermore, the cost analysis performed on realistic retail settings (e.g., Walmart-sized operations) reveals that periodic retraining can yield relevant financial savings. The estimated costs of forecasting, which scale with computing service expenses, decrease sharply when retraining is performed less frequently. This suggests that organizations can achieve economic benefits by optimizing their retraining schedules, without compromising the forecast quality.
@@
The implications of these findings are highly relevant for both academic research and industry applications. First, they challenge the prevailing assumption that forecasting models require frequent updates to maintain high predictive performance. Instead, our results suggest that global models remain effective over extended periods, meaning that retraining can be strategically planned rather than performed continuously. For practitioners, our results provide concrete guidelines on how often retraining should occur. In general, retraining every monthly appears to be a viable option to balance probabilistic accuracy and costs. Instead, if the forecasting objective is on point forecast, then even longer retraining scenarios may be adopted. Moreover, our results shed some light on the computational comparison between machine learning and deep learning models. We found that the former benefit more from less frequent retraining as the frequency of the data increases. This implies that, for large-scale applications, like the retail industry, where the forecasts of many different SKUs have to be provided regularly, machine learning models are a marginally better choice to reduce the costs of forecasting when coupled with less frequent model retraining strategies as the frequency of the data increases. These insights also have broader implications for sustainability in machine learning and AI-driven forecasting. The computational cost savings associated with reduced retraining directly translate into lower energy consumption, making forecasting operations more environmentally sustainable. This aligns with recent discussions on ”Green AI”, which emphasizes the importance of optimizing computational resources to reduce the environmental impact of machine learning applications.
@@
While our findings provide strong evidence for the feasibility of less frequent retraining in global forecasting models, some limitations remain. First, our study focused on two large retail datasets (M5 and VN1), which may not fully capture all possible time series patterns. Moreover, in this paper, we assumed that the data generating process remains stable without exhibiting significant trends or concept drifts, but in some real-world applications this assumption may simply not be valid. Future research could extend this analysis to other domains, such as financial time series or industrial production data, to verify the general validity of these findings. Furthermore, our study did not explore adaptive retraining strategies, where models are updated only when a significant drift in data distribution is detected. Such adaptive approaches could provide an even more refined balance between accuracy and computational efficiency. Future work could investigate the integration of drift detection mechanisms to optimize retraining schedules dynamically. Another direction for future studies may be the analysis of the effect of different retraining scenarios on the stability of global models, that is answering the question ”Do global models produce stable forecasts even if not continuously retrained?”. Extending the analysis to some very recent global forecasting methods that demonstrate state-of-the-art performance (e.g. transformers) would also be a natural path for future research. Finally, we encourage further studies to explore the interaction between retraining frequency and hyperparameter optimization. Indeed, different models may respond differently to retraining strategies depending on their hyperparameter configurations, and understanding these interactions could lead to even more efficient forecasting strategies.
@@
Acknowledgements
The DataLab of the University of Milano-Bicocca for the IT resources provided.
@@ Acknowledgements
LR 0.777 0.777 0.777 0.777 0.777 0.777 0.777 0.777 0.778 0.779
XGBoost 0.755 0.755 0.755 0.755 0.755 0.755 0.755 0.754 0.755 0.755
LGBM 0.771 0.772 0.771 0.772 0.772 0.772 0.772 0.772 0.772 0.772
CatBoost 0.947 0.961 0.954 0.960 0.952 0.968 0.957 0.963 0.966 0.958
MLP 0.821 0.818 0.818 0.819 0.819 0.820 0.818 0.819 0.819 0.820
TCN 0.865 0.865 0.865 0.865 0.865 0.865 0.865 0.865 0.865 0.864
NBEATSx 0.815 0.814 0.813 0.825 0.815 0.815 0.815 0.815 0.815 0.816
NHITS 0.828 0.826 0.835 0.825 0.826 0.825 0.825 0.825 0.826 0.825
@@
LR 0.267 0.268 0.269 0.269 0.270 0.271 0.272 0.271 0.272 0.274
XGBoost 0.258 0.259 0.260 0.260 0.262 0.263 0.264 0.263 0.264 0.267
LGBM 0.256 0.256 0.257 0.257 0.258 0.259 0.260 0.259 0.260 0.262
CatBoost 0.263 0.292 0.262 0.295 0.294 0.296 0.291 0.297 0.298 0.296
MLP 0.281 0.282 0.283 0.285 0.288 0.291 0.290 0.291 0.294 0.297
TCN 0.290 0.291 0.292 0.293 0.296 0.297 0.299 0.299 0.301 0.305
NBEATSx 0.279 0.280 0.281 0.288 0.286 0.289 0.289 0.289 0.292 0.295
NHITS 0.284 0.284 0.290 0.287 0.290 0.292 0.293 0.293 0.297 0.300
Table 4: M5 MQL values for each method and retrain scenario combination.
@@
LR 6.549 6.557 6.573 6.545 6.540 6.554 6.549 6.552 6.574 6.639
RF 1.868 1.869 1.869 1.865 1.864 1.862 1.858 1.859 1.848 1.835
XGBoost 1.890 1.863 1.861 1.858 1.834 1.842 1.800 1.826 1.843 1.787
LGBM 3.542 3.523 3.473 3.283 3.411 3.74 3.405 3.333 3.354 3.544
CatBoost 5.762 5.647 6.014 5.134 4.945 5.282 4.494 4.807 5.264 5.653
MLP 1.543 1.475 1.469 1.464 1.464 1.467 1.471 1.472 1.46 1.446
LSTM 1.913 1.908 1.907 1.907 1.906 1.905 1.904 1.901 1.898 1.89
TCN 1.913 1.908 1.907 1.907 1.906 1.905 1.904 1.901 1.898 1.89
NBEATSx 1.698 1.69 1.656 1.555 1.512 1.518 1.539 1.455 1.449 1.449
NHITS 1.699 1.674 1.623 1.55 1.54 1.454 1.522 1.447 1.449 1.444
@@
LR 2.896 2.916 2.953 2.949 2.949 2.953 2.959 2.999 3.043 3.136
RF 2.590 2.626 2.623 2.645 2.661 2.688 2.713 2.686 2.728 2.773
XGBoost 2.469 2.474 2.497 2.498 2.508 2.533 2.526 2.537 2.585 2.609
LGBM 2.625 2.593 2.687 2.609 2.619 2.636 2.646 2.664 2.693 2.745
CatBoost 2.845 2.823 2.936 2.728 2.736 2.813 2.739 2.728 2.774 2.860
MLP 2.492 2.410 2.420 2.406 2.414 2.432 2.430 2.463 2.471 2.503
LSTM 2.843 2.842 2.847 2.856 2.868 2.871 2.887 2.899 2.956 3.064
TCN 2.843 2.842 2.847 2.856 2.868 2.871 2.887 2.899 2.956 3.064
NBEATSx 2.626 2.613 2.580 2.456 2.470 2.507 2.495 2.450 2.480 2.568
NHITS 2.632 2.552 2.562 2.485 2.484 2.430 2.461 2.421 2.478 2.512
Table 7: VN1 MQL values for each method and retrain scenario combination.
@@
Figures 8 and 9 show the results of the Friedman-Nemenyi test on in the context of both point and probabilistic forecasting for the VN1 dataset.

### v2

@@ 4 Results and discussion
In general, the models tested performed better, in absolute terms, on the M5 dataset compared to the VN1 (see the RMSSE and SMQL tables in the Supplementary materials). This may be due to different factors like the dataset size, the frequency of the time series, the availability of external regressors to account for promotions, events, etc, and the absence of reference hyperparameter values to use during the training process.
Figure 1 shows the forecast accuracy of each model along the different retraining scenarios for the M5 and the VN1 datasets. To simplify comparisons, for each dataset, we report the results in relative terms with respect to the corresponding baseline scenario, that is $r=7$ for M5 and $r=1$ for VN1. With the exception of the CatBoost model, the RMSSE profiles are very stable over the retraining frequencies. Indeed, the accuracy remains practically the same in the M5 dataset and even improves in the VN1 dataset, regardless of the retraining scenario considered. Especially for low periodic retraining scenarios, the performance of most global models is indistinguishable from the baseline, and even if some deterioration is present for higher retraining frequencies, this is less than 5% also for the no retraining setup. These results imply that less frequent retraining does not harm the point forecast accuracy of global models. This can be explained by the fact that, if the demand patterns remain stable without significant trends or concept drifts, as is the case for both the M5 and VN1 datasets, the forecasts will accurately track demand over extended periods.
Figure 2 statistically confirms the above results for the M5 dataset 888 The Friedman-Nemenyi test is usually adopted to compare and rank the performance of different models. Here we use it to compare the accuracy produced by the same model over different retraining scenarios. . It is clear that the retraining scenarios are not statistically different in terms of point forecast accuracy at the 5% level. Indeed, even though the mean rank of the continuous retraining scenario is lower, the intervals intersect with the other scenarios, hence we can not differentiate between them. For the VN1 dataset (see Supplementary materials), less frequent retrainings are even statistically significant, meaning that periodic retraining does improve accuracy over continuous retraining.
@@
In a similar fashion to Figure 1, Figure 3 summarizes the relative accuracy in a probabilistic forecasting setting, as defined in 3.5. In this context, we observe that, for the M5 dataset, the accuracy (as measured by Multi Quantile Loss) is clearly an increasing function of the retraining scenario, meaning that less frequent updates harm the probabilistic forecasting performance of the models. This is true regardless of the method used. Nonetheless, the differences in accuracy are irrelevant for low retraining scenarios and slightly more pronounced for higher retraining levels, but in any case, less than 5-6 percentage points. We also note a small difference in the performances of machine learning and deep learning models, where the formers perform consistently better as the retraining scenario increases. For the VN1 dataset, instead, we observe an almost convex relationship between the accuracy and the retraining period. On average, models’ performance improves for low retraining scenarios and then starts deteriorating around $r=4$. The only exceptions are NBEATSx and NHITS models, designed also for long-term forecasting, which are consistently better compared to the others for longer retraining scenarios, both for point and probabilistic forecasting.
@@
The statistical test to compare probabilistic forecasting performances among the different retraining scenarios on the M5 dataset, as depicted in Figure 4, confirms that there are significant differences in accuracy. In particular, lower restraining scenarios tend to produce more accurate probabilistic forecasts, that is, less frequent retraining effectively harms the accuracy. However, this is not true for low retraining scenarios where the differences are not statistically significant, implying that even in the context of probabilistic forecasting the frequency of retraining may be reduced to at least once a month (or every two months). For the VN1 (results in Supplementary materials), instead, some level of periodic retaining is better than continuous retraining, most of the time.
@@
Figure 3: SMQL results for each method and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
Figure 4: M5 Friedman-Nemenyi test results based on SMQL.
@@
Figure 5 shows the relative computational time of each retraining scenario for the two datasets. We observe that CT decreases exponentially as the retraining scenario increases. On average, reductions are similar for both, the M5 and the VN1: going from the baseline to the first periodic retraining scenario ($r=14$ or $r=2$) almost halves CT, retraining the models every month often reduces computing time by 75%, and this reduction reaches 90% in the no retraining scenario, where the models are trained just once. However, we observe that there is a strong difference between machine learning and deep learning models in the M5 data. While the former continuously undergoes CT reductions, the latter seems to plateau at 50%, experiencing much lower gains over increasing retraining scenarios. As expected, this difference is less evident in the VN1 dataset, since it is smaller and of lower frequency compared to the M5. These results will have direct implications in terms of the costs of forecasting, which in turn may guide the choice of the model to adopt.
@@
Overall, the results of Figures 1, 3, and 5, combined with the Friedman-Nemenyi tests, suggest that retraining the global models less frequently does not harm (or even improves) forecast accuracy, while at the same time significantly reducing the computing time of producing the forecasts. Thus, extending the retraining time from the standard practice of continuous retaining to some level of periodic retraining allows for effective management of the computing time, and in turn, the costs of forecasting. Indeed, as already observed, the computational time can be directly translated into actual costs for the company. Following Nikolopoulos & Petropoulos (2018) and Fotios Petropoulos & Spiliotis (2024), we assumed some standard costs for computing services to estimate the costs of forecasting associated with each retrain scenario. Note that costs and savings are normalized by the number of series considered in each dataset (28,298 and 15,053) so that they can be directly compared and conclusions can be drawn in terms of the frequency of the time series. Figures 6, and 7 show the costs and associated savings for a large retailer, given a fixed computing service cost of \$3.5/hour, 200,000 unique SKUs and 5,000 stores, which is approximately the size of the forecasting problem of Walmart (Spiliotis & Petropoulos, 2024). As expected, the costs decrease exponentially with less frequent retraining. For daily data, forecasting with machine learning models usually costs less than with deep learning models. On average, the continuous retraining scenario costs approximately \$750,000 and this cost drops down to almost \$250,000 in the no retraining scenario, implying direct savings of more than 60%. Moreover, machine learning models allow for higher savings. Indeed, even if the models have to be updated, going from retraining every week to retraining every month allows to get direct savings of almost 75%, while it can be less than 30% for deep learning models. This implies that machine learning methods are a marginally better choice to reduce the costs of forecasting when coupled with less frequent model retraining as the frequency of the data increases. For weekly data, these differences are less evident, meaning that machine learning and deep learning models usually have very similar costs and savings profiles (without considering the Random Forest model, which is usually more than 10 times slower than other methods). This is to be expected since lower frequencies also imply smaller datasets and lower computing time. In this case, the average cost of forecasting under a continuous retraining scenario is \$250,000 (1/3 of that of daily data), and it drops to \$15,000 in the least frequent retraining scenario, producing direct savings of 90%. Nevertheless, for weekly data too, moving from retraining a model every week ($r=1$) to just once a month ($r=4$) allows to reduce the cost of almost 75%.
@@
It may be argued that for a large retailer (like Walmart), these costs (and savings) may be negligible. However, it has to be noted that the cost reduction obtained from less frequent retraining comes with no shortcomings in terms of forecasting accuracy (especially for point forecasting). Indeed, as already mentioned, models under a periodic retraining scenario result in at least the same (if not even better) forecasting performance compared to the usual practice of continuous retraining. Moreover, periodic retraining may be a good practice even when probabilistic forecasts are needed, balancing costs and accuracy effectively.
@@
Finally, in this study, since we are dealing with global models, we assumed a fixed model retraining frequency across all forecast origins and time series. While our results showed no significant difference in forecasting accuracy when reducing the retraining frequency, we believe there is potential to further improve accuracy, especially when adopting some form of periodic retraining, for instance, by allowing the retraining frequency to vary dynamically. Such adaptive strategies could involve automated procedures that monitor forecast error levels and trigger updates only when errors exceed a certain threshold. This approach would still yield computational savings, albeit smaller, since the model re-training would not be required at every review point. In this direction, Figures 8 and 9 illustrate the ”optimal” retraining frequency identified for each series in the M5 and VN1 datasets, considering all the models, based on our experiments. The results confirm that although certain series may benefit from frequent retraining, in most cases, a less frequent retraining of the models leads to better accuracy and efficiency. Indeed, for point forecasting, the optimal retraining frequency was found to be between 3 and 4 weeks (approximately 21 observations) for the M5 dataset, and around 8 to 10 weeks for the VN1 dataset. In the context of probabilistic forecasting, instead, the optimal retraining intervals were shorter (around 2 weeks for M5 and 4 weeks for VN1), suggesting that uncertainty estimation may require more frequent updates to maintain accurate levels. Moreover, while these results do not differentiate among the global forecasting methods used in the experiment, individual model differences exist, implying that even further improvements may be obtained by relying on specific model (see Supplementary material). These findings suggest that exploring intermediate retraining strategies and determining the optimal retraining frequency offers promising avenues for reducing the computational cost of the forecasting process while maintaining a high level of accuracy.
Figure 8: Overall optimal retraining frequency based on RMSSE. For each series, the optimal frequency is identified based on the average accuracy reported across the evaluation rounds per update frequency, considering all the models.
Figure 9: Overall optimal retraining frequency based on SMQL. For each series, the optimal frequency is identified based on the average accuracy reported across the evaluation rounds per update frequency, considering all the models.
@@ 5 Conclusions
In this study, we went beyond the traditional evaluation of forecasting models by also exploring the computational cost associated with generating forecasts. We examined the effects of retraining frequency on the accuracy and computational efficiency of global time series forecasting models. By systematically evaluating various retraining scenarios, ranging from continuous retraining to no retraining at all, across multiple machine learning and deep learning models, we aimed to address a key question in forecasting: is frequent retraining necessary for maintaining the high predictive performance of global models? Our findings, derived from an extensive evaluation of ten different global models across two large-scale, real-world retail datasets, challenge the conventional wisdom of continuous model retraining. Indeed, our results indicate that the common practice of continuous retraining may not always be justified, and that periodic retraining strategies can offer substantial benefits in terms of computational efficiency without a significant loss in accuracy.
@@
Our empirical evaluation provides evidence that the forecasting accuracy of global models remains stable even when the retraining frequency is significantly reduced. For point forecasting, the root mean squared scaled error (RMSSE) results indicate that models trained less frequently do not exhibit substantial performance degradation. In fact, in some cases, particularly for periodic retraining frequencies, we observe marginal improvements in accuracy. For probabilistic forecasting, the results are slightly more nuanced. While less frequent retraining does lead to minor reductions in accuracy as measured by the scaled Multi-Quantile Loss (SMQL), the degradation is relatively small (typically within 5-6%). This implies that for most practical applications, especially those where forecasting costs are key considerations, periodic retraining provides a favorable balance between accuracy and efficiency. This suggests that global models, that learn shared dynamics across multiple time series, are robust to the absence of frequent updates, especially in contexts with relatively stable demand patterns.
@@
Considering periodic retraining scenarios allows for effective management of the computing time, and in turn, the costs of forecasting. Indeed, we have shown how the computational time can be directly translated into actual costs for the company. One of the most significant findings of this study is the exponential reduction in computational costs as retraining frequency decreases. The computational time (CT) analysis demonstrates that moving from continuous retraining to a monthly retraining schedule can reduce computational costs by approximately 75%. In the extreme case of no retraining, the cost reductions approach 90%, representing a resource-saving opportunity with no shortcomings. Furthermore, the cost analysis performed on realistic retail settings (e.g., Walmart-sized operations) reveals that periodic retraining can yield relevant financial savings. The estimated costs of forecasting, which scale with computing service expenses, decrease sharply when retraining is performed less frequently. This suggests that organizations can achieve economic benefits by optimizing their retraining schedules without compromising the forecast quality.
@@
The implications of these findings are highly relevant for both academic research and industry applications. First, they challenge the prevailing assumption that forecasting models require frequent updates to maintain high predictive performance. Instead, our results suggest that global models remain effective over extended periods, meaning that retraining can be strategically planned rather than performed continuously. For practitioners, our results provide concrete guidelines on how often retraining should occur. In general, retraining every month appears to be a viable option to balance probabilistic accuracy and costs. Instead, if the forecasting objective is on point forecast, then even longer retraining scenarios may be adopted. Moreover, our results shed some light on the computational comparison between machine learning and deep learning models. We found that the former benefits more from less frequent retraining as the frequency of the data increases. This implies that, for large-scale applications, like the retail industry, where the forecasts of many different SKUs have to be provided regularly, machine learning models are a marginally better choice to reduce the costs of forecasting when coupled with less frequent model retraining strategies as the frequency of the data increases. These insights also have broader implications for sustainability in machine learning and AI-driven forecasting. The computational cost savings associated with reduced retraining directly translate into lower energy consumption, making forecasting operations more environmentally sustainable. This aligns with recent discussions on ”Green AI” which emphasizes the importance of optimizing computational resources to reduce the environmental impact of machine learning applications.
@@
While our findings provide strong evidence for the feasibility of less frequent retraining in global forecasting models, some limitations remain. First, our study focused on two large retail datasets (M5 and VN1), which may not fully capture all possible time series patterns. Moreover, in this paper, we assumed that the data-generating process remains stable without exhibiting significant trends or concept drifts, but in some real-world applications, this assumption may simply not be valid. Future research could extend this analysis to other domains, such as financial time series or industrial production data, to verify the general validity of these findings. Furthermore, our study did not explore adaptive retraining strategies, where models are updated only when a significant drift in data distribution is detected. Such adaptive approaches could provide an even more refined balance between accuracy and computational efficiency. Future work could investigate the integration of drift detection mechanisms to optimize retraining schedules dynamically. Another direction for future studies may be the analysis of the effect of different retraining scenarios on the stability of global models, that is, answering the question ”Do global models produce stable forecasts even if not continuously retrained?”. Extending the analysis to some very recent global forecasting methods that demonstrate state-of-the-art performance (e.g., transformers) would also be a natural path for future research. Finally, we encourage further studies to explore the interaction between retraining frequency and hyperparameter optimization. Indeed, different models may respond differently to retraining strategies depending on their hyperparameter configurations, and understanding these interactions could lead to even more efficient forecasting strategies.
@@
Figures 10 and 11 show the results of the Friedman-Nemenyi test in the context of both point and probabilistic forecasting for the VN1 dataset.
The costs tables 9 and 10 show the estimated cost in real values of each scenario for daily and weekly data, respectively.
Figures 12 and 13 display the optimal retraining frequency for each model in the context of the M5 dataset for point and probabilistic forecasting. Figures 14 and 15 display the optimal retraining frequency for each model in the context of the VN1 dataset for point and probabilistic forecasting. For each series, the optimal frequency is identified based on the average accuracy reported across the evaluation rounds per retraining period.
@@
LR 0.760 0.760 0.760 0.759 0.759 0.759 0.759 0.759 0.760 0.759
XGBoost 0.738 0.738 0.738 0.738 0.738 0.738 0.737 0.737 0.737 0.736
LGBM 0.754 0.754 0.754 0.754 0.754 0.754 0.754 0.754 0.754 0.754
CatBoost 0.931 0.945 0.937 0.944 0.936 0.952 0.941 0.946 0.950 0.941
MLP 0.804 0.801 0.801 0.802 0.801 0.803 0.800 0.802 0.802 0.802
TCN 0.849 0.849 0.848 0.848 0.848 0.848 0.848 0.848 0.847 0.845
NBEATSx 0.798 0.797 0.797 0.807 0.798 0.798 0.797 0.798 0.798 0.797
NHITS 0.811 0.809 0.818 0.809 0.809 0.807 0.808 0.808 0.808 0.807
@@
LR 0.268 0.268 0.269 0.269 0.271 0.272 0.273 0.272 0.273 0.274
XGBoost 0.245 0.246 0.247 0.247 0.249 0.251 0.252 0.251 0.252 0.254
LGBM 0.247 0.247 0.248 0.248 0.249 0.250 0.250 0.250 0.250 0.252
CatBoost 0.282 0.273 0.283 0.276 0.278 0.274 0.270 0.276 0.273 0.272
MLP 0.263 0.264 0.265 0.266 0.269 0.272 0.272 0.272 0.275 0.277
TCN 0.272 0.273 0.274 0.275 0.277 0.279 0.280 0.280 0.282 0.284
NBEATSx 0.262 0.263 0.264 0.269 0.268 0.271 0.271 0.271 0.274 0.276
NHITS 0.267 0.266 0.274 0.268 0.271 0.273 0.274 0.274 0.277 0.279
Table 4: M5 SMQL values for each method and retrain scenario combination.
@@
LR 7.564 7.540 7.374 7.444 7.339 7.550 7.552 7.364 7.495 7.560
RF 1.871 1.873 1.872 1.868 1.867 1.866 1.861 1.863 1.852 1.840
XGBoost 1.898 1.871 1.884 1.865 1.842 1.851 1.809 1.833 1.851 1.797
LGBM 3.540 3.520 3.471 3.282 3.414 3.743 3.447 3.355 3.382 3.561
CatBoost 6.384 6.348 6.388 5.969 5.738 6.051 5.072 5.300 5.785 6.156
MLP 1.554 1.483 1.477 1.472 1.472 1.475 1.480 1.480 1.467 1.455
LSTM 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960
TCN 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960
NBEATSx 1.704 1.696 1.661 1.561 1.519 1.525 1.544 1.462 1.455 1.456
NHITS 1.705 1.680 1.629 1.556 1.547 1.460 1.528 1.453 1.455 1.451
@@
LR 4.905 4.923 4.802 4.901 4.871 4.911 4.941 4.830 4.752 5.072
RF 1.220 1.230 1.232 1.224 1.237 1.254 1.257 1.284 1.313 1.326
XGBoost 1.503 1.496 1.493 1.505 1.483 1.488 1.466 1.451 1.515 1.479
LGBM 2.318 2.292 2.230 2.157 2.230 2.322 2.266 2.366 2.226 2.275
CatBoost 4.945 4.867 5.084 4.771 4.476 4.582 4.081 4.133 4.351 5.023
MLP 0.806 0.800 0.796 0.795 0.803 0.805 0.807 0.819 0.824 0.847
LSTM 0.892 0.895 0.899 0.903 0.910 0.917 0.924 0.935 0.957 1.003
TCN 0.891 0.895 0.899 0.903 0.910 0.917 0.924 0.935 0.957 1.003
NBEATSx 0.991 0.995 0.966 0.897 0.861 0.866 0.893 0.825 0.839 0.875
NHITS 0.992 0.996 0.944 0.894 0.876 0.814 0.880 0.820 0.833 0.862
Table 7: VN1 SMQL values for each method and retrain scenario combination.